An aeroplane taking off from the air strip has a runway of 600 m. What is its uniform acceleration if it leaves the ground in 15 seconds from the start? What is its take-off speed?
1
Expert's answer
2019-10-28T10:57:54-0400

The distance traveled by an aeroplane before it taking off the ground


"d=\frac{at^2}{2}"

So, the acceleration


"a=\frac{2d}{t^2}=\frac{2\times 600}{15^2}=5.3\:\rm m/s^2"

The take-off speed


"v=at=5.3\times 15=80\:\rm m/s"